OpenStudy (anonymous):

a company is selling a new book in two sizes the dimensions of the first book is are 28inches long and the 12 inches wide. if the ratio or scale factor between the two books is 4. find the dimensions of the larger book?

OpenStudy (anonymous):

you are going to take 28x4 and 12x4 because a scale factor is like multiplying

OpenStudy (anonymous):

Isn't it the square root of the scale factor that you multiply into the length and widths? area of first book = 28 * 12 = 336. area of larger book = 4 * 336 = 1344. 2*28 = 56 inches long, 2*12 = 24 inches wide. 56/24 = 28/12 and 56*24 = 1344.

OpenStudy (anonymous):

Thanks everyone i really appreciate it. I am in the 7th grade and this stuff is hard thanks again